# Hierarchical Light Transformer Ensembles for Multimodal Trajectory Forecasting

### v1

@@ 6 Conclusion
In conclusion, our proposed loss and ensembling framework, HLT-Ens, introduces a novel and efficient approach to multimodal trajectory forecasting, achieving state-of-the-art under identical experimental conditions. HLT-Ens leverages a hierarchical density representation and associated loss to train Mixture Density Networks and benefits from efficient ensemble training inspired by grouped convolutions.
In light of these advancements, HLT-Ens emerges as a promising approach with the capacity to enhance the performance and robustness of multimodal trajectory forecasting systems. This innovation opens avenues for improved decision-making in autonomous systems and related domains, marking a significant step forward in the evolution of trajectory prediction methodologies.
Appendix 0.A Notations
@@
$\boldsymbol{\pi}_{\theta}$ The probability vector corresponding to the mixture weights
$\overline{\boldsymbol{\mu}}_{\theta_{k^{\star}}}$ The mean of the $k$th meta-mode of a Laplace mixture distribution parametrized by $\theta$
$\overline{\mathbf{b}}_{\theta_{k^{\star}}}$ The scale vector of the $k$th meta-mode of a Laplace mixture distribution parametrized by $\theta$
@@
Appendix 0.B Implementation and Training Details
@@
$K^{\prime}$ Number of modes within a meta-mode for our approaches 3 3 3 3
@@
This section details both the implementation and training procedure of our models concerning the experiments. Upon acceptance and with the validation of our industrial sponsor, we expect to release the corresponding code. We implement all networks using the PyTorch framework and leverage a Nvidia RTX A6000 to train them. For both datasets, we used the same simple preprocessing:
@@
Appendix 0.C HWTA Loss Details
@@
This section provides more explanations and insights on the HWTA loss. First, we present how classification loss terms are encompassed within the meta-mode loss and the MWTA loss. Then we illustrate our loss behavior on a 2D toy dataset (Fig. 4).
@@
$\mathcal{L}^{\mathcal{C}}_{meta}$ $=\frac{1}{K^{\star}}\sum_{Z^{\star}}D_{\text{KL}}(Q(Z^{\star})||P_{\theta}(Z^{\star}|Y,X))$ (23)
@@
where $Z^{\star}$ and $Z^{\prime}$ are discrete latent variables corresponding to the meta-modes and their respective modes. $D_{\text{KL}}(.||.)$ is the Kullback-Leibler divergence between the approximated posterior and the actual posterior. As in [13], we set $Q(Z^{\star})=P_{\theta_{\text{old}}}(Z^{\star}|Y,X)$ and $Q(Z^{\prime})=P_{\theta_{\text{old}}}(Z^{\prime}|Y,X)$, where $\theta_{\text{old}}$ are the parameters before the optimization step. Eq. (7) becomes:
@@
and their respective probabilities being $P(S_{1})=P(S_{4})=\frac{1-t}{2}$, $P(S_{2})=P(S_{3})=\frac{t}{2}$, and $P(S_{5})=0$. Whenever a region is selected, a point is sampled from it uniformly. Fig. 4(a) illustrates such distribution for $t\in\{0,0.5,1\}$. To have a better illustration of the loss dynamics, we train a basic three-layer fully connected network with $50$ neurons in each hidden layer and ReLU activation function, similar to [35]. Given the time $t$, we are interested in modeling a two-dimensional distribution using $K=10$ modes. We visually compare the effect of the mixture NLL loss, the WTA loss combined with the NLL loss, and our HWTA loss (assuming $2$ meta-modes here) in Fig. 4(b). Although the NLL loss seems to capture better the underlying distribution, the mode diversity is compromised by some redundant modes, diminishing their coverage. The WTA loss on the mixture NLL creates a Voronoï tesselation of the space, i.e., the modes are efficiently placed for coverage. Yet, it provides no information on what mode to keep if we were to reduce the number of modes subsequently. On the contrary, the hierarchy in our loss enables us to give $2$ levels of modeling. To maximize coverage and diminish the overall number of modes, one can directly take the meta-modes. We illustrate the benefit of this strategy in Appendix 0.F.
Appendix 0.D Loss Parameter Sensitivity Study
@@
This section showcases the effect of $\gamma$ on our HWTA loss $\mathcal{L}$. In particular, Fig. 5 illustrates the variation of performance for different $\gamma$ values. Note that for $\gamma=0.0$ the meta-mode loss $\mathcal{L}_{\text{meta}}$ is not used and for $\gamma=1.0$ the loss $\mathcal{L}_{\text{MWTA}}$ is annealed. From our experiments, $\mathcal{L}_{\text{meta}}$ is necessary for better accuracy on most confident predictions as we observe a dramatic decrease in $\text{mADE}_{1}$ and $\text{mFDE}_{1}$
Appendix 0.E Importance of the Width Factor
@@
Our efficient ensembling architecture depends on two hyperparameters. $M$ corresponds to the ensemble size, and $\alpha$, as a factor on the embedding size, controls the width of the DNN. We evaluate the sensitivity of HLT-Ens to these parameters by training models using ADAPT backbone on the Interaction dataset with various settings. Tab. 6 and Tab. 7 showcase the effect of $\alpha$ for $4$ and $8$ subnetworks respectively. Increasing $\alpha$ enables better results until it reaches a plateau at a cost of more parameters.
Appendix 0.F Robustness Analysis over Mode Number
@@
The number of modes is a critical hyperparameter significantly impacting the model’s performance. To cover the true multimodal distribution it is often preferred to predict more modes and choose subsequently the most confident ones. Yet, doing so might decrease the diversity in the predicted modes as the most confident mode is likely to have duplicates. Using hierarchy in the mixture distribution, we showcase that we are able to reduce our mixture complexity while retaining most of its diversity. Fig. 6 illustrates the impact of the number of modes over the performance on diversity metrics (i.e., $\text{mADE}_{6}$ and $\text{mFDE}_{6}$). The comparison is done with three different settings (Legacy loss, WTA loss, and HWTA loss) all using AutoBots backbone and trained on Argoverse 1. Concerning our loss, instead of taking the $6$ most confident trajectories, we directly use the $6$ meta-modes as a set of forecasts. Doing so seems to stabilize the performance, suggesting our loss enables more robustness to variations in the number of modes with a fixed number of meta-modes.
Appendix 0.G Diversity in Ensembles of Mixtures
@@
Figure 7: Similarity matrices for DE and HLT-Ens on Argoverse 1. We use the AutoBots backbone, to construct ensembles of size $M=3$. For HLT-Ens, we set $\alpha=1.5$.
Prediction diversity is essential for the performance of ensembles. In [22], the authors present two sources of stochasticity in the training process producing diversity among ensemble members: the random initialization of the model’s parameters and the shuffling of the batches. HLT-Ens does not benefit from this last source of stochasticity, yet has comparable performance. To provide more insight into the effect on the diversity of this source of stochasticity in the training of trajectory forecasting models, we conducted a small experiment on the consistency of the cluster found by the post-hoc KMeans algorithm on our ensembles.Fig. 7 presents the similarity matrices for both DE and HLT-Ens. Each cell represents the rate of two modes being clusterized in the same cluster by the KMeans algorithm on the validation set of Argoverse 1. It appears that the clustering is more consistent for HLT-Ens highlighting the possibility to cluster only once the modes and apply it after without too much performance loss compared to executing a KMeans algorithm for each sample.
Appendix 0.H SceneTransformer Experiments
Table 8: Performance comparison on Interaction using SceneTransformer backbone. All ensembles have $M=3$ subnetworks and are followed by a KMeans algorithm to form $6$ trajectory clusters from which we take the centroids; we highlight the best performances in bold. For our method, we consider $\alpha=1.5$. The number of parameters is expressed in millions. M.A. corresponds to the inference cost, i.e., the number of Mega multiply-add operations for a forward pass, which is estimated with the python package torchinfo [39].
Method $\text{mADE}_{1}$downarrow $\text{mADE}_{3}$downarrow $\text{mADE}_{6}$downarrow $\text{mFDE}_{1}$downarrow $\text{mFDE}_{3}$downarrow $\text{mFDE}_{6}$downarrow $\text{NLL}_{3}$downarrow $\text{NLL}_{6}$downarrow #Prm $\downarrow$
Single Single model
Legacy 0.66 0.31 0.21 1.50 0.66 0.39 18.57 17.23
HWTA (Ours) 0.48 0.32 0.26 1.17 0.71 0.52 -4.74 -11.15 11.8
@@ Appendix 0.H SceneTransformer Experiments
DE 0.57 0.28 0.21 1.30 0.60 0.39 23.09 22.59 35.4
HT-Ens (Ours) 0.50 0.32 0.23 1.19 0.69 0.47 -7.34 -9.94 35.4
HLT-Ens (Ours) 0.49 0.32 0.25 1.17 0.72 0.53 -4.24 -9.59 8.9
Tab. 8 presents the performance of our method on the SceneTransformer backbone trained on the Interaction dataset. These results are based on a custom re-implementation of SceneTransformer we developed. These preliminary results showcase the ability of our approach to improve the most confident forecast accuracy as well as the quality of the predicted multimodal distribution compared to the original loss of SceneTransformer. However, it seems to lack diversity in the predicted modes. We argue it might be necessary to try out other values of $\gamma$ (we used $\gamma=0.6$ here).

### v3

@@ 5 Experiments
Qualitative Results:
Fig. 4 illustrates our new loss behavior. As expected, the sub-modes of the same meta-mode are near each other. This highlights the ability of our approach to enable fast compression of the mixture distribution by taking the meta-modes only. Doing so would retain the most diversity while simplifying the forecast. We explore this aspect in Appendix F and provide more qualitative results in Appendix I.
@@ 6 Conclusion
In conclusion, our proposed HWTA loss and ensembling framework, HLT-Ens, introduces a novel and efficient approach to multimodal trajectory forecasting, achieving state-of-the-art under identical experimental conditions. HLT-Ens leverages a hierarchical density representation and associated loss to train Mixture Density Networks and benefits from efficient ensemble training inspired by grouped convolutions.
In light of these advancements, HLT-Ens emerges as a promising approach to structuring better multi-modal distributions, improving the performance and robustness of multi-modal trajectory forecasting systems. Yet, the computation cost of aggregation strategies for ensembles remains an obstacle to their industrial applications. We believe spatial information provided by hierarchical multi-modal predictions and parallel ensemble training could help find more efficient ways to merge forecasts.
7 Acknowledgment
This work was granted access to the HPC resources of IDRIS under the allocation 2023-AD011014689 made by GENCI.
Appendix A Notations
@@
$\boldsymbol{\pi}_{\theta}$ The probability vector corresponding to the a mixture weights
$\bar{\boldsymbol{\mu}}_{\theta_{k}}$ The mean of the $k$th meta-mode of a Laplace mixture distribution parametrized by $\theta$
$\bar{\mathbf{b}}_{\theta_{k}}$ The scale vector of the $k$th meta-mode of a Laplace mixture distribution parametrized by $\theta$
@@
Table 4: Summary of the main notations of the paper.
Appendix B Implementation and Training Details
@@
$K^{\prime}$ Number of modes within a meta-mode for our approaches 2 3 3 3
@@
This section details our models’ implementation and training procedure for the experiments. We expect to release the corresponding code upon acceptance and validation by our industrial sponsor. We implement all networks using the PyTorch framework and train them using an Nvidia RTX A6000, except for experiments on Wayformer where a Nvidia Tesla V100 has been used. For both datasets, we used the same simple preprocessing:
@@
Appendix C HWTA Loss Details
@@
This section provides more explanations and insights on the HWTA loss. First, we present how classification loss terms are encompassed within the meta-mode and MWTA losses. Then, we illustrate our loss behavior on a 2D toy dataset (Fig. 5).
@@
$\mathcal{L}^{\mathcal{C}}_{meta}$ $=\frac{1}{K}\sum_{Z}D_{\text{KL}}(Q(Z)||P_{\theta}(Z|Y,X))$ (23)
@@
where $Z$ and $Z^{\prime}$ are discrete latent variables corresponding to the meta-modes and their respective modes. $D_{\text{KL}}(.||.)$ is the Kullback-Leibler divergence between the approximated posterior and the actual posterior. As in [14], we set $Q(Z)=P_{\theta_{\text{old}}}(Z|Y,X)$ and $Q(Z^{\prime})=P_{\theta_{\text{old}}}(Z^{\prime}|Y,X)$, where $\theta_{\text{old}}$ are the parameters before the optimization step. Eq. (7) becomes:
@@
and their respective probabilities being $P(S_{1})=P(S_{4})=\frac{1-t}{2}$, $P(S_{2})=P(S_{3})=\frac{t}{2}$, and $P(S_{5})=0$. Whenever a region is selected, a point is sampled from it uniformly. Fig. 5(a) illustrates such distribution for $t\in\{0,0.5,1\}$. To better illustrate the loss dynamics, we train a basic three-layer fully connected network with $50$ neurons in each hidden layer and ReLU activation function, similar to [37]. Given the time $t$, we are interested in modeling a two-dimensional distribution using $K=10$ modes. We visually compare the effect of the mixture NLL loss, the WTA loss combined with the NLL loss, and our HWTA loss (assuming $2$ meta-modes here) in Fig. 5(b). Although the NLL loss seems to capture the underlying distribution better, the mode diversity is compromised by some redundant modes, diminishing their coverage. The WTA loss on the mixture NLL creates a Voronoï tesselation of the space, i.e., the modes are efficiently placed for coverage. Yet, it provides no information on what mode to keep if we subsequently reduce the number of modes. On the contrary, the hierarchy in our loss enables us to give $2$ levels of modeling. One can directly take the meta-modes to maximize coverage and diminish the overall number of modes. We illustrate the benefit of this strategy in Appendix F.
Appendix D Loss Parameter Sensitivity Study
@@
This section showcases the effect of $\gamma$ on our HWTA loss $\mathcal{L}$. In particular, Fig. 6 illustrates the performance variation for different $\gamma$ values. Note that for $\gamma=0.0$ the meta-mode loss $\mathcal{L}_{\text{meta}}$ is not used and for $\gamma=1.0$ the loss $\mathcal{L}_{\text{MWTA}}$ is annealed. From our experiments, $\mathcal{L}_{\text{meta}}$ is necessary for better accuracy on most confident predictions as we observe a dramatic decrease in $\text{mADE}_{1}$ and $\text{mFDE}_{1}$
Appendix E Importance of the Width Factor
@@
Our efficient ensembling architecture depends on two hyperparameters. $M$ corresponds to the ensemble size, and $\alpha$, as a factor on the embedding size, controls the width of the DNN. We evaluate the sensitivity of HLT-Ens to these parameters by training models using the ADAPT backbone on the Interaction dataset with various settings. Tab. 7 and Tab. 8 showcase the effect of $\alpha$ for $4$ and $8$ subnetworks respectively. Increasing $\alpha$ enables better results until it reaches a plateau at the cost of more parameters.
Appendix F Robustness Analysis over Mode Number
@@
The number of modes is a critical hyperparameter significantly impacting the model’s performance. Predicting more modes and choosing the most confident ones is often preferred to cover the true multimodal distribution. Yet, doing so might decrease the diversity in the predicted modes as the most confident mode is likely to have duplicates. Using hierarchy in the mixture distribution, we showcase that we can reduce our mixture complexity while retaining most of its diversity. Fig. 7 illustrates the impact of the number of modes over the performance on diversity metrics (i.e., $\text{mADE}_{6}$ and $\text{mFDE}_{6}$). The comparison is done with three different settings (Legacy loss, WTA loss, and HWTA loss), all using AutoBots backbone and trained on Argoverse 1. Concerning our loss, instead of taking the $6$ most confident trajectories, we use the $6$ meta-modes as a set of forecasts. Doing so seems to stabilize the performance, suggesting our loss enables more robustness to variations in the number of modes with a fixed number of meta-modes.
Appendix G Diversity in Ensembles of Mixtures
@@
Figure 8: Similarity matrices for DE and HLT-Ens on Argoverse 1. We use the AutoBots backbone to construct ensembles of size $M=3$. For HLT-Ens, we set $\alpha=1.5$.
Prediction diversity is essential for the performance of ensembles. In [23], the authors present two sources of stochasticity in the training process producing diversity among ensemble members: the random initialization of the model’s parameters and the shuffling of the batches. HLT-Ens does not benefit from this last source of stochasticity, yet it has comparable performance. To provide more insight into the effect on the diversity of this source of stochasticity in the training of trajectory forecasting models, we conducted a small experiment on the consistency of the cluster found by the post-hoc KMeans algorithm on our ensembles. Fig. 8 presents the similarity matrices for both DE and HLT-Ens. Each cell represents the rate of two modes being clusterized in the same cluster by the KMeans algorithm on the validation set of Argoverse 1. The clustering appears more consistent for HLT-Ens (i.e., the similarity matrix is sparser), highlighting the possibility of clustering only once the modes and applying it without too much performance loss compared to executing a KMeans algorithm for each sample.
Appendix H Wayformer and SceneTransformer Experiments
Method $\text{mADE}_{1}$downarrow $\text{mADE}_{6}$downarrow $\text{mFDE}_{1}$downarrow $\text{mFDE}_{6}$downarrow $\text{NLL}_{3}$downarrow $\text{NLL}_{6}$downarrow #Prm $\downarrow$
SceneTransformer Single model
Legacy 0.66 0.21 1.50 0.39 18.57 17.23
HWTA (Ours) 0.48 0.26 1.17 0.52 -4.74 -11.15 11.8
@@ Appendix H Wayformer and SceneTransformer Experiments
DE 0.57 0.21 1.30 0.60 23.09 22.59 35.4
HT-Ens (Ours) 0.50 0.23 1.19 0.47 -7.34 -9.94 35.4
HLT-Ens (Ours) 0.49 0.25 1.17 0.53 -4.24 -9.59 8.9
Wayformer Single model
WTA 1.75 0.51 4.13 1.20 49.92 -11.58
HWTA (Ours) 0.81 0.40 2.24 0.95 -15.54 -49.71 1.1
Ensemble
DE 1.01 0.48 2.74 1.19 -12.81 -21.23 3.3
HT-Ens (Ours) 0.74 0.35 2.06 0.80 -25.80 -47.36 3.3
HLT-Ens (Ours) 0.68 0.33 1.87 0.71 -27.64 -49.78 1.5
Table 9: Performance comparison on Interaction for the SceneTransformer and Wayformer backbones. All ensembles have $M=3$ subnetworks and are followed by a KMeans algorithm to form $6$ trajectory clusters from which we take the centroids; we highlight the best performances in bold. For our method, we consider $\alpha=1.5$ for SceneTransformer and $\alpha=2.0$ for Wayformer. The number of parameters is expressed in millions.
Tab. 9 presents the performance of our method on the SceneTransformer and Wayformer backbones trained on the Interaction dataset. These results are based on custom re-implementations we developed for both methods. These preliminary results showcase the ability of our approach to improve the most confident forecast accuracy and the quality of the predicted multimodal distribution compared to the original loss presented in both papers. Our method applied to Wayformer outperforms its counterpart on all metrics. Interestingly, HLT-Ens even outperforms HT-Ens, making it an attractive choice. It only has slightly more parameters than the classic Wayformer. Concerning SceneTransformer, we observe a lack of diversity in our predicted modes. We argue it might be necessary to try out other values of $\gamma$ (we used $\gamma=0.6$ here).
Appendix I Additional Qualitative Results
Fig. 9 provides additional trajectory forecasting examples on the Argoverse 1 dataset. We display predictions from an AutoBots trained using our new loss HWTA and a classical AutoBots model. The first thing we observe is that the predictions from our model are less scattered than the other. Indeed, they are closer to each other and the ground truth, which explains why we reached higher performance on $\text{mADE}_{1}$, $\text{mFDE}_{1}$ and the $\text{NLL}_{k}$ metrics. We also note that the sub-modes belonging to the same meta-modes are near each other, as announced in the paper. With this knowledge, we can expect the average of several sub-modes (i.e., a meta-mode) to be a more robust prediction, as it might be in higher-density areas. Finally, we show a failure example, where due to the bad position of one sub-mode, the meta-mode A is misplaced. We argue that this issue could occur for likely trajectory candidates and that one could easily compute the intra-mode distances to see whether the cluster is coherent. Moreover, adding more sub-modes per meta-mode should alleviate this issue as one bad sub-mode will have less effect.
(a) Left turn scenario
(b) Intersection scenario
(c) Another intersection scenario
Figure 9: Qualitative results with AutoBots backbone on Argoverse 1. We compare an AutoBots model trained with its original loss ($K=6$) compared with our HWTA loss ($\gamma=0.8$) with $3$ meta-modes (i.e., $K^{\star}=3$ and $K^{\prime}=2$).